# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

### v1

@@ 5 Discussion
Now, we delve into another regime, characterized by a relatively large energy of the vector field. In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small. Many questions need to be reconsidered in this regime, such as the anisotropy for one vector field and non-Gaussianity. We posit that the discussions in this paper are also applicable to a single vector field. In this case, the $IN^{2}$ term in power spectrum, which served as a contribution to anisotropy for small $h$ case, is replaced by other constant modes. It is imperative for us to calculate the anisotropy in this regime. Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration. Regarding PBHs, we are discussing the isotropic configuration of vector fields as an example. Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking. These differences manifest in the generated PBHs and induced gravitational waves. We intend to present these studies in the near future.
@@ Appendix A Derivation of the quadratic action
Appendix B Growing solution of the adiabatic mode
@@
$\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (B.2)
and $\mu\equiv p/H$ and $u\equiv m_{s}^{2}/H^{2}+32h^{2}$. It’s useful to define the variable $y$ as $u^{2}y^{2}=u^{2}+128h^{2}\mu^{2}$ Then the integration in (B.1) becomes
$\mathcal{I}(N)\equiv\frac{u}{8\sqrt{2}h}\int_{y_{1}}^{y_{2}}\sqrt{b\left(y-1\right)-\left(y^{2}-1\right)}\frac{ydy}{y^{2}-1}$ (B.3)
@@
$y_{1}=\sqrt{1+\frac{128h^{2}\mu^{2}}{u^{2}}},\ \ \ \ \ \ \ y_{2}=\sqrt{1+\frac{128h^{2}|m_{s}|^{2}}{u^{2}H^{2}}}.$ (B.4)
@@
$-\frac{\sqrt{2b}}{2}\arctan\left[\frac{2(1-y)+b(y-1-2)}{2\sqrt{2b}\sqrt{1-b+by-y^{2}}}\right],$ (B.5)
@@
$-\frac{\sqrt{2b}}{2}\left[\frac{\pi}{2}-\arctan\left[\frac{2(1-y_{1})+b(y_{1}-3)}{2\sqrt{2b}\sqrt{1-b+by_{1}-y_{1}^{2}}}\right]\right]\Bigg{\}}.$ (B.7)
@@
$u^{2}\simeq\frac{8h^{2}(1+6h^{2})}{1+2h^{2}},\ \ \ \ \ \ \ b\simeq\frac{8(1+2h^{2})}{1+6h^{2}},\ \ \ \ \ \ \ y_{1}\simeq 1.$ (B.8)
@@
$\mathcal{I}_{\text{approx}}=\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (B.9)
@@
Appendix C Bogoliubov coefficients
@@
$h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (C.1)
@@
^$\hat{\mathcal{F}}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (C.2)
@@
^$\hat{\mathcal{F}}^{\text{I}}(t_{1}^{-})=$$\hat{\mathcal{F}}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{I}}(t_{1}^{-})=\dot{\hat{\mathcal{F}}}^{\text{II}}(t_{1}^{+}),$ (C.3)
@@
^$\hat{\mathcal{F}}^{\text{II}}(t_{2}^{-})=$$\hat{\mathcal{F}}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{II}}(t_{2}^{-})=\dot{\hat{\mathcal{F}}}^{\text{III}}(t_{2}^{+}).$ (C.5)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i+\kappa-S_{\pm}\right)\Bigg{\}},$ (C.6)
@@
$H_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots.$ (C.7)

### v2

@@ 5 Discussion
Now, we delve into another regime, characterized by a relatively large energy of the vector field. In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small. Many questions need to be reconsidered in this regime, such as the anisotropy for one vector field and non-Gaussianity. We posit that the discussions in this paper are also applicable to a single vector field. In this case, the $IN^{2}$ term in power spectrum, which served as a contribution to anisotropy for small $h$ case, is replaced by other constant modes. It is imperative for us to calculate the anisotropy in this regime. Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration. Regarding PBHs, we are discussing the isotropic configuration of vector fields as an example. Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking. These differences manifest in the generated PBHs and induced gravitational waves. Recently observations from 15-year NANOGrav data set reveals information about low-frequency stochastic gravitational wave background [67]. PBHs can be one of the sources of this background [68]. The formation of PBHs in single-field up to one-loop correction is still controversial [69, 70, 71, 72, 73, 74, 75, 76]. The single-field EFT with a imaginary $c_{s}$ should also be discussed. We intend to present these studies in the near future.
@@ Appendix A Derivation of the quadratic action
Appendix B Effective field theory of inflation
$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (B.1)
$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (B.2)
$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (B.3)
In the low-energy scale $\omega^{2}\ll p^{2}+M^{2}$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity, the Lagrangian of NG boson is given by [40]
$\mathcal{L}_{\pi}=$ $\sqrt{-g}\Bigg{\{}-M_{\text{pl}}^{2}\dot{H}\left[\dot{\pi}\left(1+\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}|\dot{H}|}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\pi)^{2}\right]$
$+2M_{2}^{4}\left[\dot{\pi}^{2}-(\tilde{\partial}\pi)^{2}\right]\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\dot{\pi}-\frac{4}{3}M_{3}^{4}\left(\dot{\pi}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)^{2}\dot{\pi}+\cdots\Bigg{\}},$ (B.4)
where we have expanding the action up to third order of fluctuations.
$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}\epsilon\left[\dot{\pi}\left(\frac{M^{2}c_{s}^{-2}-\tilde{\partial}^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\mathcal{\pi})^{2}\right],$ (B.5)
where the speed of sound is defined by
$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (B.6)
The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model. The dispersion relation of the propagation of the light field $\pi$ should be written as
$\omega^{2}(p)=\frac{M^{2}+p^{2}}{M^{2}c_{s}^{-2}+p^{2}}p^{2},$ (B.7)
which is generally non-linear. However, if the speed of sound $c_{s}$ is not very small, The region with nonlinear dispersion relations is very narrow so the Lagrangian of NG boson shortly reduces to
$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (B.8)
where the dispersion relation is almost linear $\omega(p)\simeq c_{s}p$ throughout the entire process.
Appendix C Growing solution of the adiabatic mode
@@
$\frac{|\omega_{-}|}{H}=\sqrt{-\mu^{2}-\frac{1}{2}u+\frac{1}{2}\sqrt{u^{2}+128h^{2}\mu^{2}}}$ (C.2)
and $\mu\equiv p/H$ and $u\equiv m_{s}^{2}/H^{2}+32h^{2}$. It’s useful to define the variable $y$ as $u^{2}y^{2}=u^{2}+128h^{2}\mu^{2}$ Then the integration in (C.1) becomes
$\mathcal{I}(N)\equiv\frac{u}{8\sqrt{2}h}\int_{y_{1}}^{y_{2}}\sqrt{b\left(y-1\right)-\left(y^{2}-1\right)}\frac{ydy}{y^{2}-1}$ (C.3)
@@
$y_{1}=\sqrt{1+\frac{128h^{2}\mu^{2}}{u^{2}}},\ \ \ \ \ \ \ y_{2}=\sqrt{1+\frac{128h^{2}|m_{s}|^{2}}{u^{2}H^{2}}}.$ (C.4)
@@
$-\frac{\sqrt{2b}}{2}\arctan\left[\frac{2(1-y)+b(y-1-2)}{2\sqrt{2b}\sqrt{1-b+by-y^{2}}}\right],$ (C.5)
@@
$-\frac{\sqrt{2b}}{2}\left[\frac{\pi}{2}-\arctan\left[\frac{2(1-y_{1})+b(y_{1}-3)}{2\sqrt{2b}\sqrt{1-b+by_{1}-y_{1}^{2}}}\right]\right]\Bigg{\}}.$ (C.7)
@@
$u^{2}\simeq\frac{8h^{2}(1+6h^{2})}{1+2h^{2}},\ \ \ \ \ \ \ b\simeq\frac{8(1+2h^{2})}{1+6h^{2}},\ \ \ \ \ \ \ y_{1}\simeq 1.$ (C.8)
@@
$\mathcal{I}_{\text{approx}}=\sqrt{2}\pi\left(2-\sqrt{\frac{1+6h^{2}}{1+2h^{2}}}\right)h.$ (C.9)
@@
Appendix D Bogoliubov coefficients
@@
$h(N)=h_{0}\left[\theta(t-t_{1})-\theta(t-t_{2})\right].$ (D.1)
@@
^$\hat{\mathcal{F}}^{\text{I}}(t,\boldsymbol{k})=$ $u_{k}(t)\hat{a}_{\mathcal{F}}(\boldsymbol{k})+\text{h.c.}(-\boldsymbol{k}),$ (D.2)
@@
^$\hat{\mathcal{F}}^{\text{I}}(t_{1}^{-})=$$\hat{\mathcal{F}}^{\text{II}}(t_{1}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{I}}(t_{1}^{-})=\dot{\hat{\mathcal{F}}}^{\text{II}}(t_{1}^{+}),$ (D.3)
@@
^$\hat{\mathcal{F}}^{\text{II}}(t_{2}^{-})=$$\hat{\mathcal{F}}^{\text{III}}(t_{2}^{+}),\ \ \ \ \ \ \ \dot{\hat{\mathcal{F}}}^{\text{II}}(t_{2}^{-})=\dot{\hat{\mathcal{F}}}^{\text{III}}(t_{2}^{+}).$ (D.5)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -e^{-iS_{\pm}\delta N}\left[\kappa^{2}+S_{\pm}\left(i+\kappa\right)\right]\left(i+\kappa-S_{\pm}\right)\Bigg{\}},$ (D.6)
@@
$H_{\zeta}=$ $\frac{e^{-3\delta N/2}}{3\sqrt{3}}\left(2e^{3\delta N}-1-i\frac{6}{x_{2}^{3}}-ie^{2\delta N}\frac{3}{x_{2}}\right)\sin{\left(2\sqrt{6}h\delta N\right)}+\cdots.$ (D.7)